Find a quadratic model for the set of values
(-2,8),(0,-4),(4,68)

Answer by richwmiller(17219)

quadratic equation model
f(x)=ax^2+bx+c
plug in the values of to find ab and c
f(-2)=8
a(-2)^2+b(-2)+c=8
4a-2b-4=8
4a-2b=12
f(0)=-4
a(0)+b*0+c=-4
c=-4
f(4)=68
a(4)^2+b(4)-4=68
16a+4b=72
4a-2b=12
multiply the second by 2
16a+4b=72
8a-4b=24
add them
24a=96
a=4
4(4)-2b=12
16-2b=12
4=2b
2=b
a=4 b=2 c=-4
y=4x^2+2x-4 quadratic equation
y=2(2x^2+x-2)